OpenStudy (jazzyluvsu123):

i need to know if these are right??? 12 - 3& 2/9 = 8/3 4- 1& 5/8 = 37/8 5 - 3& 1/2 = 5 did I do this right or is it like this--- 11& 9/9 - 3& 2/9 = 8& 7/9 @Euler271 @agent0smith @ranga

OpenStudy (ranga):

No they are incorrect. Do you want to show the steps you took for the first one. You get the correct way to do one you can do the rest.

OpenStudy (jazzyluvsu123):

ok yes plz show me i added on to my question did u see wat i added on

OpenStudy (ranga):

"or is it like this--- 11& 9/9 - 3& 2/9 = 8& 7/9" That is correct!

OpenStudy (agent0smith):

12 - 3& 2/9 i showed how to do in the last question

OpenStudy (jazzyluvsu123):

yes i think this is how you did it 11 9/9 - 3 2/9 = 8 7/9 right

OpenStudy (ranga):

12 can be written as: 11 + 9/9 3&2/9 = 3 + 2/9 12 - 3&2/9 = (11 + 9/9) - (3 + 2/9) = 11 + 9/9 - 3 - 2/9 = 11 - 3 + 9/9 - 2/9 = 8 +(9 - 2) / 9 = 8 + 7/9 = 8&7/9

OpenStudy (jazzyluvsu123):

thanks for helping i got it now

OpenStudy (ranga):

cool. you are welcome.